X has a normal distribution with a mean of 80.0 and standard deviation of 3.5. Find the following probabilites:
a) P(x<77.0)
b) P(78.0<0x<85.0)
c) P(x>90.0)

Answer by stanbon(75887)

X has a normal distribution with a mean of 80.0 and standard deviation of 3.5. Find the following probabilites:
---
Comment: For each of these problems find the z-value.
Then use your calculator or z-chart to find the appropriate
probability.
----
a) P(x<77.0)
z(77) = (77-80)/3.5 = -0.8571
P(x < 77) = P(z < -0.8571) = 0.1957
=======================================
Follow the same procedure for "b" and "c".
Cheers,
Stan H.
b) P(78.0<0x<85.0)
c) P(x>90.0)